OpenStudy (anonymous):

What is the value of the expression 4!(7 – 4)! i need help i dont understand this problem.. A. 144 B. 5040 C. 120,384 D. 479,001,600

OpenStudy (anonymous):

First, note that ! means factorial. That means 2! = 2 x 1. 3! = 3 x 2 x 1. This wants you to find 4! and multiply that by (7 - 4) factorial. We can write it like this:\[(4 \times 3 \times 2 \times 1)(7 -4)!\] We can simplify this to: \[24(3)!\] That means 24 (3 x 2 x 1), which is 144, the answer is A.

OpenStudy (anonymous):

ok great i just got it thanks!